If the ratio of sums of first p terms and q terms of an arithmetic progressio is p 2 :q 2 then the ratio of its pth and qth term is? 2p-1:2q-1 2p+1:2q+1 p:q p+1:q+1

  1. If the ratio of sums of first p terms and q terms of an arithmetic progressio is p2:q2 then the ratio of its pth and qth term is?
  1. 2p-1:2q-1
  2. 2p+1:2q+1
  3. p:q
  4. p+1:q+1

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Sum of p terms,
Sp = p/2 [2a + (p-1)d]
Sum of q terms,
Sq = q/2[2a + (q-1)d]
And given 
Sp/Sq = p^2/q^2
do it and sort out further u would get somthing something relation in pth and qth terms relation.